OpenStudy (anonymous):

Find the area of the largest isosceles triangle that can be inscribed in a circle of radius 6 inches.

jhonyy9 (jhonyy9):

if the radius is 6 inches than the height of this triangle is 9 inches because the radius is 2/3 from this height so than h=(6*3)/2=9

jhonyy9 (jhonyy9):

area equal base multiply height divide 2 = (b*9)/2

OpenStudy (anonymous):

 all angles are 60 degrees. 2/3h=r 2/3h=6 h=9 then since you have a right triangle you can have the base as (9tan30)=sqrtof3 over 3 then solve for area of triangle 1/2 bh = 1/2(sqrt3 over 3)(9) but this is just 1 triangle so you need to multiply it by 2 so just bh = sqrt of 3 over 3 * 9

OpenStudy (anonymous):

approximately 5.2 squared inches